# CrystalFormer-RL: Reinforcement Fine-Tuning for Materials Design

### v2

@@ Abstract
Reinforcement fine-tuning played an instrumental role in enhancing the instruction-following and reasoning abilities of large language models. In this work, we employ reinforcement fine-tuning for materials design, in which discriminative machine learning models are used to provide rewards to the autoregressive transformer-based materials generative model CrystalFormer [1]. By optimizing the reward signals—such as energy above the convex hull and material property figures of merit—reinforcement fine-tuning infuses knowledge from discriminative models into generative models. The resulting model, CrystalFormer-RL, shows enhanced stability in generated crystals and successfully discovers crystals with desirable yet conflicting material properties, such as substantial dielectric constant and band gap simultaneously. Notably, we observe that reinforcement fine-tuning not only enables the property-guided material design but also unlocks property-based material retrieval behavior of pretrained generative model. The present framework opens an exciting gateway to the synergies of the machine learning ecosystem for materials design.
@@ III.2 Reinforcement learning from the figure of merits for property-guided material generation
Table 1: Dielectric materials discovered by CrystalFormer-RL, ranked according to their figures of merit (FoM): $\varepsilon_{\mathrm{elec}}\cdot E_{g}$. The band gap, dielectric constant, and energy above the convex hull are all calculated by DFT. The two highlighted rows are not in the Alex-20 dataset. See details of the samples at https://drive.google.com/file/d/1bEGzeDPybiDZw5lCzLwRapoMSbaGfQN0/view?usp=sharing.
@@ Appendix C Details of reinforcement fine-tuning implementation
$\mathcal{L}^{\mathrm{PPO}}(\theta)=\mathop{\mathbb{E}}_{x\sim p_{\mathrm{old}}(x)}\left[\mathrm{min}(\gamma(\theta)A(x),\mathrm{clip}(\gamma(\theta),1-\epsilon,1+\epsilon))A(x)\right],$ (S1)
where $\gamma(\theta)=\frac{p_{\theta}(x)}{p_{\mathrm{old}}(x)}$ denotes the probability ratio between the policy under optimization and the policy for sampling, $A(x)$ is the advantage function, and $\epsilon$ is the clipping parameter. The second term, $\mathrm{clip}(\gamma(\theta),1-\epsilon,1+\epsilon)$, modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $\gamma(\theta)$ outside of the interval $1-\epsilon,1+\epsilon]$. Drawing samples from $p_{\mathrm{old}}(x)$ rather than $p_{\theta}(x)$ stems from the importance sampling and reduce the number of costly evaluation of the reward function.
We use $R(x)$ denotes the reward function with entropy regularization,
$R(x)=r(x)-\tau\ln\frac{p_{\theta}(x)}{p_{\mathrm{base}}(x)}.$ (S2)
The advantage function $A(x)$ is then defined as
A(x)=R(x)b,A(x)=R(x)-b, (S3)
where $b$ represents the baseline. Note that one does not need to compute the gradient of $A(x)$ with respect to network parameters $\theta$ in the PPO algorithm. Instead, one uses the score function gradient estimate [79] to maximize the objective function $\mathcal{L}^{\mathrm{PPO}}(\theta)$. For each batch of samples from $p_{\mathrm{old}}$ and reward signal, we carry out a few gradient ascent steps to update the policy network. The detailed hyperparameters used in the PPO algorithm are listed in Table S1.
The advantage function measures the deviation of the reward $R(x)$ from this baseline, indicating how much better or worse the generated samples is compared to the baseline. In standard PPO implementations, this baseline is typically approximated by a value function, which is trained alongside the policy model. We use the average reward of the generated samples $x$, i.e., $b=\mathop{\mathbb{E}}_{x\sim p_{\mathrm{old}}}[R(x)]$ as the baseline [80]. This approach obviates the need for training a separate value function and is equivalent to the group relative policy optimization done right [67, 81].

### v3

# Reinforcement Fine-Tuning for Materials Design

@@ Abstract
Reinforcement fine-tuning played an instrumental role in enhancing the instruction-following and reasoning abilities of large language models. In this work, we employ reinforcement fine-tuning for materials design, in which discriminative machine learning models are used to provide rewards to the autoregressive transformer-based materials generative model CrystalFormer [Sci. Bull. 70, 3522 (2025)]. By optimizing the reward signals—such as energy above the convex hull and material property figures of merit—reinforcement fine-tuning infuses knowledge from discriminative models into generative models. The resulting model, CrystalFormer-RL, shows enhanced stability in generated crystals and successfully discovers crystals with desirable yet conflicting material properties, such as substantial dielectric constant and band gap simultaneously. Notably, we observe that reinforcement fine-tuning not only enables the property-guided material design but also unlocks property-based material retrieval behavior of pretrained generative model. The present framework opens an exciting gateway to the synergies of the machine learning ecosystem for materials design.
@@ III.2 Reinforcement learning from the figure of merits for property-guided material generation
Table 1: Dielectric materials discovered by CrystalFormer-RL, ranked according to their figures of merit (FoM): $\varepsilon_{\mathrm{elec}}\cdot E_{g}$. The band gap, dielectric constant, and energy above the convex hull are all calculated by DFT. The two highlighted rows are not in the Alex-20 dataset. See details of the samples at https://zenodo.org/records/17982949.
@@ Appendix C Details of reinforcement fine-tuning implementation
$\mathcal{L}^{\mathrm{PPO}}(\theta)=\mathop{\mathbb{E}}_{x\sim p_{\mathrm{old}}(x)}\left[\mathrm{min}(\gamma(\theta)A(x),\mathrm{clip}(\gamma(\theta),1-\epsilon,1+\epsilon)A(x))\right],$ (S1)
where $\gamma(\theta)=\frac{p_{\theta}(x)}{p_{\mathrm{old}}(x)}$ denotes the probability ratio between the policy under optimization and the policy for sampling, $A(x)$ is the advantage function, and $\epsilon$ is the clipping parameter. The second term, $\mathrm{clip}(\gamma(\theta),1-\epsilon,1+\epsilon)$, modifies the surrogate objective by clipping the probability ratio, which removes the incentive for moving $\gamma(\theta)$ outside of the interval $1-\epsilon,1+\epsilon]$. Drawing samples from $p_{\mathrm{old}}(x)$ rather than $p_{\theta}(x)$ enables sample reuse and allows multiple gradient update steps on the same batch of data, thereby improving sample efficiency without requiring costly re-evaluation of the reward function for each update.
The advantage function $\mathcal{A}(x)$ is then defined as
$\mathcal{A}(x)=r(x)-b,$ (S2)
where $b$ represents the baseline. For simplicity, we incorporate the KL regularization term into the advantage function, yielding
$A(x)=\mathcal{A}(x)-\tau\ln\frac{p_{\theta}(x)}{p_{\mathrm{base}}(x)}.$ (S3)
Note that one does not need to compute the gradient of $A(x)$ with respect to network parameters $\theta$ in the PPO algorithm. Instead, one uses the score function gradient estimate [79] to maximize the objective function $\mathcal{L}^{\mathrm{PPO}}(\theta)$. For each batch of samples from $p_{\mathrm{old}}$ and reward signal, we carry out a few gradient ascent steps to update the policy network. The detailed hyperparameters used in the PPO algorithm are listed in Table S1.
The advantage function measures the deviation of the reward $r(x)$ from this baseline, indicating how much better or worse the generated samples is compared to the baseline. In standard PPO implementations, this baseline is typically approximated by a value function, which is trained alongside the policy model. We use the average reward of the generated samples $x$, i.e., $b=\mathop{\mathbb{E}}_{x\sim p_{\mathrm{old}}}[r(x)]$ as the baseline [80]. This approach obviates the need for training a separate value function and is closely related to the group relative policy optimization done right [67, 81].